250 g of 50% HNO₃ were diluted with 150 g of water. Calculate the mass percentage of diluted acid.


How do I solve this? Step 1 to finish please.


1
Expert's answer
2020-10-28T08:41:43-0400

"250g" initial solution contains "50\% \,\, HNO_3"


"\therefore" mass of "HNO_3" = "= 50\% \times 250g = \dfrac{50}{100} \times 250g = 125g"



Mass of final solution = "250g" of "50\% \, HNO_3" solution + "150g" of Water = "450g"




Mass percentage of the acid =

"\dfrac{mass \ of \ acid}{mass \ of \ final \ solution} \times 100\%"


"= \dfrac{125}{400} \times 100\%"


"= 31.25 \%"